Write the expression as the sine or cosine of an angle.
sin ( pi/2 ) cos (pi/7) + cos ( pi/2) sin (pi/7)

Answer by ikleyn(53213)

.

sin (pi/2) = 1;   cos (pi/2) = 0.  Therefore,


sin ( pi/2 ) cos (pi/7) + cos ( pi/2) sin (pi/7) = 1*cos(pi/7) + 0*sin(pi/7) = cos(pi/7).


Solved, explained and completed.